Sarah is twice the age Linda is. 6 years ago, Sarah was four times the age of Linda. What was their ages six years ago as well as their ages now?
Side note: M'techer sad it ain't 12, 3, 18, and 9.

Answer by LinnW(1048)

Set x = Sarah's current age
Set y = Linda's current age
Sarah's age 6 years ago was (x-6)
Linda's age 6 years ago was (y-6)
x = 2y
(x-6)=4(y-6)
Now substitute 2y for x in the equation above
((2y)-6)=4(y-6)
2y-6 = 4y - 24
add -4y to each side
-2y - 6 = -24
add 6 to each side
-2y = -18
divide each side by -2
y = 9
Since x = 2y, x = 18
Sarah is currently 18 and Linda is 9
Six years ago they were 12 and 3
*****************
I am not sure what M'techer is thinking.